### v2

# Li Lv Wang Luo Zhang Schuller

Abstract
Speech deepfake detection (DFD) has benefited from diverse acoustic and semantic speech representations, many of which encode valuable speech information and are costly to train. Existing approaches typically enhance DFD by tuning the representations or applying post-hoc classification on frozen features, limiting control over improving discriminative DF cues without distorting original semantics. We find that emotion is encoded across diverse speech features and correlates with DFD. Therefore, we introduce a unified, feature-agnostic, and non-destructive training framework that uses emotion as a bridging constraint to guide speech features toward DFD, treating emotion recognition as a representation alignment objective rather than an auxiliary task, while preserving the original semantic information. Experiments on FakeOrReal and IntheWild show accuracy improvements of up to 6% and 2%, respectively, with corresponding reductions in equal error rate. Code is in the supplementary material.
keywords:
Speech, Deepfake detection, Emotion, Feature selection, Pretraining
1 Introduction
Speech deepfake detection (DFD) in audio heavily relies on the selection of appropriate and robust feature representations [iqbal2022deepfake]. Physical acoustic features are among the most commonly used, including Mel Frequency Cepstral Coefficients (MFCCs) [hamza2022deepfake], chroma features [bakken2025deep], and comprehensive sets of acoustic descriptors extracted using tools such as openSMILE [pascu2025easy, eyben2010opensmile]. In parallel, deep learning–based features derived from pretrained models such as Whisper [radford2023robust], regarded as raw audio physical features, have also shown promising results. Beyond these low-level representations, high-level task-oriented features have proven effective as well. These are often obtained from models trained for specific downstream applications, such as automatic speech recognition (ASR) [mansoor2024audio] and speaker verification (SV) [pianese2022deepfake], and tend to offer more representations with semantic information compared to simple physical features.
Given the extracted features, a common approach is to train the entire network end-to-end by jointly optimising the feature extractors and the downstream classifier. Under this setting, pretrained models such as Whisper or wav2vec 2.0 [luo2024whisper, wang2025awaveformer] are directly fine-tuned for DFD, resulting in substantial structural adaptation of the original representations toward DFD features. While effective for detection, full fine-tuning substantially alters the feature space and limits the reuse of these representations in future tasks. Preserving their semantic information is therefore crucial, as the same features may later be required after DFD for tasks such as ASR and SV in multimodal large language models [xu2025qwen2]. Moreover, end-to-end fine-tuning typically needs to be repeated for each new deepfake dataset, as the learned representations may overfit to dataset-specific characteristics [khanjani2023audio]. On the other hand, to reduce training cost, an alternative post-hoc strategy directly applies a classifier head on top of frozen speech representations for DFD [saha2024exploring]. While this approach is computationally efficient, it treats the extracted features as fixed and provides no structured mechanism to guide the representations toward deepfake-discriminative characteristics, often resulting in limited performance gains.
Despite these advances, the field of audio DFD remains fragmented, with no unified training framework that both preserves the semantic information of the original features and guides them toward deepfake-discriminative representations.
Notably, we observe that the emotion recognition task provides a shared factor that can guide heterogeneous speech representations toward deepfake discrimination. Firstly, emotion can be expressed through a variety of feature sets, including low-level descriptors of physical characteristics [kishore2013emotion, li2025gatedxlstmmultimodalaffectivecomputing] as well as higher-level application-based features derived from ASR systems [tits2018asr]. For instance, the Macro-Voice framework disentangles speaker identity from emotional cues to obtain a more purified emotion representation, highlighting the intersection between speaker verification representations and emotional features [tian2025marcovoicetechnicalreport]. These observations suggest that most acoustic features capture emotional information and can be used for emotion recognition. Moreover, emotion cues have been shown to provide effective discriminative information for audio DFD [conti2022deepfake, mittal2020emotions] and are inherently difficult for generative models to reproduce authentically over successive iterations, making them a robust and reliable signal for DFD in practical applications [sun2024friendlyaicomprehensivereview, mobbs2025emotion, li2025artificialemotionsurveytheories]. Therefore, emotion recognition can act as a bridge for leveraging heterogeneous speech features in DFD.
To address the lack of a unified, feature-preserving approach for audio DFD, we make the following contribution. We propose a training pipeline that introduces a pre-training stage prior to the final post-hoc classifier, without fine-tuning the feature extractors and requiring only a lightweight classifier for detection. In the pre-training stage, we found emotion, as a suitable feature-agnostic constraint applicable to arbitrary speech representations, guiding them toward deepfake-discriminative cues while preserving their original semantic structure. This design enables deepfake-oriented representation adjustment with lower training cost than full fine-tuning.
Our method differs fundamentally from prior studies that demonstrate the effectiveness of emotion features in DFD by aligning emotion with other modalities [mittal2020emotions, hosler2021deepfakes]. These works primarily focus on incorporating additional modalities, such as visual information, to improve emotion recognition, and directly employ emotion representations as input features for detection. Moreover, our method differs from conventional fusion strategies, such as early or late fusion, that combine emotion features with other audio representations [lei2025deepfake, alsaeedi2025audio]. These approaches rely on explicit feature concatenation or selection, which requires careful feature engineering and exhibits limited scalability as feature diversity increases. In contrast, we do not treat emotion as a standalone feature nor perform explicit feature fusion; instead, we leverage emotion as a bridging constraint to guide heterogeneous representations toward deepfake discrimination without replacing their original semantic structures. Experimental results show consistent improvements on the FakeOrReal (FoR) [reimao2019dataset] and InTheWild (ITW) [muller2022does] benchmarks, specifically up to approximately 6% and 2% increases for accuracy, respectively, and in equal error rate (EER), showing reductions of up to about 4% and 1%, and comparable performance on ASVSpoof2019 LA [wang2020asvspoof].
2 Methodology: Emotion-Guided Representation Alignment (EmoBridge)
To address the aforementioned challenge of unifying a training pipeline for preserving speech representation to enhance DFD, we align affective cues through an emotion prediction head that jointly optimises existing features with an emotion-related task. This emotion constraint adaptation enables the feature representation to incorporate new affective information without forgetting previously learned knowledge. The overall framework is illustrated in Figure 1.
Figure 1: Overview of the proposed Emotion-Guided Representation Alignment (EmoBridge) framework for speech DFD. In Phase I, raw speech is transformed into speech representations through two parallel routes: hand-crafted acoustic feature extractors (e.g., OpenSMILE) and encoders pre-trained for other speech tasks such as ASR or speaker verification. For hand-crafted acoustic features, a subset of features containing richer emotion-related cues is retained. For deep representations, an emotion prediction head is attached to the pre-trained encoder and optimized using an emotion recognition objective, guiding the representation to incorporate emotion-related information while preserving its original task-oriented structure. In Phase II, the resulting emotion-aligned representations—either selected acoustic features or outputs from the trained encoder—are frozen and fed into a downstream classifier (e.g., SVM or MLP) to perform DFD, classifying input speech as real or fake.
Given the vast variety of feature representations, we consider the last hidden states of the feature encoder (for attention-based architectures) or the final layer outputs (for other deep learning architectures) as the optimal feature representations for each modality. We categorise commonly used features into three groups: (1) acoustic features, which can be extracted using openSMILE; (2) deep learning (DL) raw features, derived directly from deep neural models; and (3) application-based features, such as semantic features from ASR pre-trained models or perceptual features from SV pre-trained models.
In detail, to investigate the effect of emotion-guided representation alignment, we select four representative types of speech features: hand-crafted acoustic features extracted using openSMILE [eyben2010opensmile]; low-level deep learning features from Whisper [radford2023robust] [^1] https://huggingface.co/openai/whisper-small ; semantic and linguistic features from the ASR model SpeechT5 [ao2022speecht5unifiedmodalencoderdecoderpretraining] [^2] https://huggingface.co/microsoft/speecht5_asr ; and perceptual features from the speaker verification model WavLM [Chen_2022] [^3] https://huggingface.co/microsoft/wavlm-base-sv. These pre-trained models are widely adopted in their respective domains, and we maintain comparable model configurations across experiments to ensure fair and consistent evaluation.
For hand-crafted features, we utilise openSMILE tools to directly extract acoustic features (e. g., F0, MFCCs). For this model, our analysis focuses on comparing two functional-level feature sets: eGeMAPSv02, a set designed by experts and IS09, a feature set from a challenge, both are designed for emotion recognition. The latter emphasizes more emotional cues in addition to basic acoustic features. For other feature sets, we attach an emotion prediction head implemented as a multilayer perceptron (MLP) to the feature encoder and jointly optimize the encoder and the emotion head on an emotion recognition task using a general-purpose emotion dataset. During the EmoBridge stage, encoder parameters are updated through the emotion objective only, without retaining the original pretraining objective. To construct a diverse and representative training corpus, we combine four widely used emotion recognition datasets: the Toronto Emotional Speech Set (TESS) [dupuis_pichora-fuller_2010_tess], Surrey Audio-Visual Expressed Emotion (SAVEE) [jackson_haq_savee], CREMA-D [cao2014crema-d], and RAVDESS [livingstone2018ravdess]. All audio from these datasets is used to train the EmoBridge stage to avoid emotion dataset bias as much as possible, while we use Emotion Speech Dataset (ESD) [zhou2021esd] for testing the learned emotion representations.
After completing the EmoBridge stage, the emotion-aligned encoder is connected to a downstream classifier to perform the final speech DFD task. To evaluate the effectiveness of the emotion-guided representations, we extract the outputs from each encoder layer and feed them individually into a support vector machine (SVM) [cortes1995svm] for classification. During this stage, the encoder is kept frozen to avoid feature interference and to ensure that DFD performance solely reflects the quality of the learned representations. The average performance across all encoder layers is reported to quantify the overall contribution of the proposed training framework. Our approach is scalable and feature-agnostic. The selected experimental features span a broad range of acoustic, linguistic, and perceptual representations, and serve as representative examples to demonstrate the general applicability of the EmoBridge framework, which can be readily extended to other speech feature extractors. When no suitable pre-trained model is available for a given feature set, a corresponding encoder can be trained during a pre-training stage and subsequently incorporated into the proposed pipeline. Accordingly, the proposed framework can be formally expressed as Equation 1.
$\begin{aligned} \theta^{*}_{pre}&=\arg\max_{\theta}\log P(y_{A}\mid x_{A};\theta),\text{ feature pretraining}\\ \theta^{*}_{emo+pre},\phi^{*}_{emo}&=\arg\max_{\theta,\phi}\log P(y_{B}\mid x_{B};\theta^{*}_{pre},\phi),\text{ emotion alignment}\\ \varsigma^{*}_{df}&=\arg\max_{\varsigma}\log P(y_{C}\mid x_{C};\theta^{*}_{emo+pre},\varsigma),\text{ DFD}\end{aligned}$
where $\theta$ denotes the parameters of model where feature information is stored. $\phi$ is the parameter of the emotion prediction head, and $\varsigma$ is the parameter of of the DFD classifier. $(x_{A},y_{A})$, $(x_{B},y_{B})$, and $(x_{C},y_{C})$ are the input-output pairs for speech feature extraction (e.g. ASR, SV), emotion recognition and DFD tasks, respectively, and $\mathcal{D}_{A}$, $\mathcal{D}_{B}$, and $\mathcal{D}_{C}$ are their respective data distributions. $P(y\mid x;\cdot)$ represents the conditional probability of the label given the input, which corresponds to the predictive distribution used in negative log-likelihood.
3 Experiments and result
3.1 Dataset and experiments setup
We select the ASVSpoof2019 LA subset [wang2020asvspoof] and the FoR [reimao2019dataset] and ITW [muller2022does] datasets as benchmarks for the DFD task. For ASVSpoof2019 LA and ITW, we follow the official dataset splits in the original papers. For FoR, we use the for-norm data splits. This setup results in around 71k, 4.5k, and 32k test samples for ASVSpoof2019 LA, FoR, and ITW, respectively.
We employ a three-layer fully connected network with hidden dimensions of 768, 256, and 7 to train the emotion recognition task. The model is optimized using AdamW with a learning rate of 1e-5 for a maximum of 40 epochs. To ensure a fair comparison, we also evaluate a pre-trained model that already incorporates emotion-related features, namely HuBERT [hsu2021hubertselfsupervisedspeechrepresentation] pre-trained on the IEMOCAP dataset [busso2008iemocap] [^4] https://huggingface.co/superb/hubert-base-superb-er. The emotion recognition results are summarised in Table 1. The models achieve commendable results on the emotion test sets, indicating that the emotion constraining process is successful. Although they do not reach the state-of-the-art models for emotion recognition, this is expected, as they retain the original features and not overfited on emotion recognition.
Table 1: EmoBridge performance. The table reports the weighted accuracy for the four pre-trained models evaluated in this study, with comparison to other models in literature.
Metrics ASR SV DL-raw Emotion I. Shahin [shahin2019emotion] S. Hamsa [hamsa2020emotion] Koya S. [koya2022ea]
Weighted Acc. 0.728 0.701 0.786 0.732 0.840 0.909 0.917
After completing the EmoBridge step, we apply the default settings of SVM from the scikit-learn library [^5] https://scikit-learn.org/stable/api/sklearn.svm.html to each layer of the trained EmoBridge models, extracting both low-level and high-level feature representations, which better represent the model feature fusion. Model performance is then evaluated using average accuracy and EER as metrics.
3.2 Results
Table 2 reports the layer-wise averaged EER and accuracy of the four models and acoustic features, evaluated on the FoR, ITW, and ASVSpoof2019 LA datasets. This evaluation protocol reflects the robustness of the representation across different abstraction levels and avoids bias to particular layers.
Table 2: Comparison of layer-wise averaged EER and accuracy between the Pre-trained model and the model with the emotion as a bridge strategy on three datasets.
Dataset Model Pre-trained EmoBridge
EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
FoR openSMILE .406 .574 .275 .663
Emotion .139 .854 .078 .913
SV .136 .858 .089 .908
ASR .082 .887 .045 .946
DL-raw .091 .890 .062 .931
ITW openSMILE .200 .897 .188 .897
Emotion .050 .953 .047 .953
SV .061 .942 .058 .946
ASR .054 .948 .052 .947
DL-raw .033 .969 .020 .982
ASVSpoof 2019 LA openSMILE .406 .574 .275 .663
Emotion .061 .970 .071 .966
SV .070 .964 .079 .960
ASR .076 .960 .078 .960
DL-raw .052 .946 .049 .967
This demonstrates that incorporating the EmoBridge strategy consistently improves the model’s discriminative capability for deepfake compared to the pre-trained baseline across most datasets and feature types. Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy. Compared with the Emotion-pretrained HuBERT model, the performance before EmoBridge training is not dominant. We attribute this to the fact that HuBERT is trained on IEMOCAP, which provides a limited coverage of emotional expressions. By incorporating more diverse emotion knowledge, the detection performance improves noticeably, indicating that a broader and balanced emotion distribution is necessary.
Moreover, across all datasets, the DL-raw features consistently benefit from the EmoBridge, indicating that raw deep learning representations capture key cues for DFD and that guidance provided by emotion information is effective. The FoR corpus benefits the most from our strategy, showing substantial reductions in EER (e. g., from 0.082 to 0.045 for ASR features) along with notable accuracy gains. Improvements on the ITW dataset are also observed. Although the relative gains are modest, the large scale of the dataset means that these gains translate into a considerable number of additional correctly classified samples. These findings suggest that emotion-based bridging provides the greatest advantage in diverse or emotionally expressive conditions.
In contrast, comparable performance is observed on ASVSpoof. To identify where our approach is most vulnerable, we further analyse model performance across different sources within the test set. We also compare against two existing emotion-based methods: using the static emotion feature [conti2022deepfake] (results reported directly from the original paper), and a concatenating emotion feature with others (adapted from [mittal2020emotions]), which we combine HuBERT embeddings with DL-raw features (using our same experimental configuration of the models in Section 3.1). The results are shown in Table 3.
Table 3: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy. Column Static is from work [mittal2020emotions] (— means not provided or N/A) while Column Emotion is from our strategy using Hubert as original feature. In this table, only the most noteworthy results are emphasised in bold.
Source ASR SV DL-raw Emotion Static Concatenation
EmoBridge? No Yes No Yes No Yes No Yes
A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998 0.948 0.994
A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999 0.988 0.962
A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977 0.900 0.965
A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999 0.895 0.997
A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998 0.890 1.000
A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.831 1.000
A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000 0.763 0.952
A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999 0.927 0.954
A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993 0.898 0.988
A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984 0.988
A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975 0.869
A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903 0.905
Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564 0.832
The results indicate that the relatively smaller gains observed on ASVSpoof2019 LA can be attributed to the highly imbalanced distribution of bona fide and spoofed samples in this dataset, where real speech constitutes only a small portion of the training data. Nevertheless, the DL-raw results still achieve higher performance on the real subset and improved overall accuracy, indicating that emotion-enhanced representations remain beneficial. Audio generated by methods A17–A19 shows more consistent improvements than other source types, likely because these methods are from real human speech, which contains richer prosodic and emotional variation than TTS-based synthesis. This observation does not suggest a limitation of the proposed framework, but highlights the importance of bona fide diversity when emotion is used as an alignment signal.
3.3 Effects of DFD classifier
We select SVM as the main classifier for simplicity, although other classifiers could also be used. Previous work shows that the relative performance of different feature representations remains consistent across classifier choices [chakravarty2024lightweight], even though absolute scores may vary. To further validate the robustness of our approach, we conduct an ablation study on the FoR dataset using a two-layer MLP classifier (hidden dimension 512 with a 2-dimensional output), as shown in Table 4. The results show similar performance across feature types, demonstrating the robustness of our method.
Table 4: EER and accuracy between the Pre-trained model and the model with the EmoBridge strategy on FoR with MLP as classifier.
Dataset Model Pre-trained Emotion fused
EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
FoR openSMILE .385 .554 .265 .645
Emotion .143 .866 .083 .915
SV .146 .864 .087 .910
ASR .088 .898 .066 .952
DL-raw .095 .899 .077 .943
3.4 Preservation of speech representation
Additionally, we evaluated the preservation of speech representation via feature visualization using selected samples from the EmoFake [zhao2024emofake]. These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations. We analyze SV and ASR features, as shown in Figure 2, using t-SNE [maaten2008visualizing] for visualization.
After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion guidance. Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, although the clusters are slightly less compact than before EmoBridge training. Directly evaluating ASR performance via word error rate after EmoBridge is non-trivial, as ASR systems rely on a tightly coupled encoder–decoder architecture, and re-training or re-adapting the decoder would introduce additional confounding factors. We therefore assess representation preservation through encoder-level, task-agnostic evaluation. Specifically, we compare speaker verification performance in a zero-shot setting. Using the same settings as in [aldeneh2024remove], the models before and after EmoBridge achieve EERs of 0.313 and 0.325, respectively, on VoxCeleb [nagrani2020voxceleb], indicating no huge degradation. This suggests that the proposed strategy avoids fully altering the original semantic information and prevents overfiting to a specific downstream task, enhancing scalability.
Figure 2: t-SNE for SV (left) and ASR (right) visualizations of selected sample representations obtained before (model_ori) and after (model_new) Emobrigde. The x_y_zx\_y\_z label denotes the speaker, emotion, and content respectively; identical subscripts indicate the same speaker, emotion, or content.
3.5 Explainability for model improvement
Our approach yields performance gains, so we further investigate the model’s internal mechanism. Specifically, we examine the mean attention values across time steps for different layers before and after applying our strategy, using the DL-raw features as an example, shown in Figure 3. The results show that the model after applying our strategy retains partial overlap with the original attention distribution, indicating that it preserves useful prior information, while also developing new attention regions where more affective cues are likely captured.
Figure 3: Comparison of mean attention values of Whisper layers before (model_ori) and after emotion as a bridge (model_new).
4 Conclusion
We propose EmoBridge, a unified strategy that leverages emotion information as a constraint to guide pre-trained feature embeddings. This approach preserves the original feature representations while injecting affective cues, enabling the model to better expose discriminative patterns for DFD and consistently outperform baseline methods across multiple benchmarks. Our experiments show that emotion information provides the effective guidance for improving representation quality in this framework. Future work could explore integrating multimodal features with emotion-aware guidance for DFD.
5 Generative AI Use Disclosure
We only use Generative AI to check the grammar issue of our paper.